I need to translate this problem into an equation........
there are twelve people on the jury. THere are 4 more men than women. How many men are there?

Answer by oberobic(2304)

Let x be the number of women on the jury. So x+4 = number of men.
x + (x+4) = 12.
2x + 4 = 12
2x = 8
x = 4
So the number of women is 4.
The number of men is 8.
The number of men is 4 more than the number of women, so the answer checks.